OpenStudy (anonymous):

solve by factoring the following quadratic 4x^2-x^2-4x+1

OpenStudy (anonymous):

To solve (find the roots of) the quadratic equation =0: First, we'll attempt to factor this quadratic equation as a trinomial. 4x^2-x^2-4x+1which is equal to 3x^2-4x+1 is a valid trinomial that might be factorable. This trinomial can be factored. Here is it in factored form: 3x^2-4x+1= ( x-1 )( 3x-1 ) With the two factors found, we now solve the equations: x-1 =0 and 3x-1 =0. Solving x-1 =0, we get that x= 1 Solving 3x-1 =0 and, we get that x=1/3 Therefore, x= 1 and x=1/3 are the roots to the quadratic equation 3x^2-4x+1